# DeepSeek vs. ChatGPT: A Comparative Study for Scientific Computing and Scientific Machine Learning Tasks

### v1

@@ Abstract
Large Language Models (LLMs) have emerged as powerful tools for tackling a wide range of problems, including those in scientific computing, particularly in solving partial differential equations (PDEs). However, different models exhibit distinct strengths and preferences, resulting in varying levels of performance. In this paper, we compare the capabilities of the most advanced LLMs—ChatGPT and DeepSeek—along with their reasoning-optimized versions in addressing computational challenges. Specifically, we evaluate their proficiency in solving traditional numerical problems in scientific computing as well as leveraging scientific machine learning techniques for PDE-based problems. We designed all our experiments so that a non-trivial decision is required, e.g. defining the proper space of input functions for neural operator learning. Our findings reveal that the latest model, ChatGPT o3-mini-high, usually delivers the most accurate results while also responding significantly faster than its reasoning counterpart, DeepSeek R1. This enhanced speed and accuracy make ChatGPT o3-mini-high a more practical and efficient choice for diverse computational tasks at this juncture.
@@ 1 Introduction
Large language models (LLMs), such as ChatGPT [1, 2, 3] and, most recently, DeepSeek [4] are general-purpose artificial intelligence (AI) chatbots based on the transformer architecture [5], trained on large amounts of data from the Internet, and have learned to engage in conversations based on user requests. Among their many applications, there has been growing interest in leveraging LLMs for scientific research, particularly in coding, mathematics, and problem-solving. This interest has driven rapid recent advancements in LLM development. For example, on January 31, 2025, OpenAI released OpenAI o3-mini [2], the latest reasoning model supposed to deliver particular strength in science, math, and coding. This came only a few days after DeepSeek unveiled its own reasoning model DeepSeek-R1 [6] that also specializes in math, coding, and reasoning tasks. These models build on prior advancements such as ChatGPT-4o [1] and DeepSeek V3 [4], marking a new phase of competition among LLM developers seeking to optimize performance in scientific fields.
@@
With the rising prominence of DeepSeek and the intensifying competition among LLM developers, particularly between DeepSeek and OpenAI, this study conducts a comparative testing of their models in computational mathematics and scientific machine learning. Deliberately tricky problems are formulated and posed to both reasoning and non-reasoning models from DeepSeek and OpenAI. We first test the LLMs’ capability in numerical methods, including numerical integration, finite difference methods (FDM), and finite element methods (FEM). Since machine learning methods have gained popularity as alternatives to traditional numerical approaches [36, 37], we also assess the LLMs’ performance in scientific machine learning tasks, including image recognition, physics-informed learning with physics-informed neural networks (PINNs) [38], and operator learning methods such as Deep Operator Network (DeepONet) [39]. By comparing the models’ performances and their reasoning decisions across a variety of problems, the work provides insight into how well DeepSeek and OpenAI models perform in computational mathematics.
@@ 2 Experiments
In this section, we evaluate DeepSeek and OpenAI LLMs by testing them on a range of challenging benchmark problems in numerical algorithms and scientific machine learning. The models under consideration are:
@@
    ChatGPT 4o [1]: OpenAI’s flagship model designed for general reasoning and multimodal capabilities, including text, code, and vision processing. It employs an optimized transformer architecture and focuses on improving efficiency and reducing latency while maintaining performance in coding, mathematics, and logical inference.
@@
    ChatGPT o3-mini-high [2]: ChatGPT o3-mini-high is a variant of OpenAI’s o3-mini model, optimized for tasks requiring intensive reasoning in coding, mathematics, and science. Despite its smaller size, o3-mini-high is supposed to deliver performance comparable to larger models by employing advanced reasoning techniques.
To ensure fairness, all chat memory and user-personalized settings are disabled. This prevents models from benefiting from prior context and ensures that each query is treated independently. Additionally, since LLM response time can be significantly affected by server latency rather than the LLM’s evaluation speed, we do not compare or report response times. Instead, our evaluation focuses on the quality of the generated solutions and, for reasoning models, the reasoning process and decision-making. To differentiate the model performances, the test problems selected are tricky but also leave enough flexibility for the LLMs to make their own decisions in what method and parameters to use in their solutions. Most problems are advanced, typically at the PhD level, and demand a deep understanding of computational mathematics and familiarity with recent research advancements. The coding languages tested are chosen to be TensorFlow [40] and Jax [41] in Python, which are frameworks frequently used in the scientific machine learning community.
@@ 2.1 Traditional numerical methods
We present the four LLMs under consideration with a variety of problems in numerical methods and solving differential equations. We compare the performance of their generated solutions and discuss the decisions the LLMs made.
@@ 2.1.1 Numerical solution to stiff ODEs
The Robertson Problem Equation 1 is a famous example for a stiff system of ODEs that describes the chemical reaction of H. H. Robertson [42].
@@
The LLMs’ responses are summarized in Table 1, and the computed solutions are plotted in Figure 1. For reference, the ODE system is also solved in scipy using an implicit Runge-Kutta method of Radau IIA family of order 5 (Radau) [43]. The RK4 schemes implemented by DeepSeek V3 and ChatGPT 4o failed to solve the problem and yielded exponentially growing solutions, while the backward Euler schemes implemented by the reasoning models DeepSeek R1 and ChatGPT o3-mini-high both gave the correct solution. Additionally, ChatGPT o3-mini-high implements an adaptive time stepping that compares one full step of size $\Delta t$ with two half‑steps (each of size $\Delta t/2$) to estimate the local error. If the error is below a prescribed tolerance, the step is accepted; otherwise, it is rejected and a new (smaller) time step is computed.
@@
DeepSeek V3 DeepSeek R1 ChatGPT 4o ChatGPT o3-mini-high
Method Fourth Order Backward Euler Fourth Order Backward Euler with
Runge-Kutta (RK4) Runge-Kutta (RK4) adaptive time stepping
Step size $0.01$ $0.1$ $0.1$ adaptive with
initial step size $10^{-6}$
$L_{2}$ Error (%) $x$ NaN $7.86\times 10^{-6}\%$ NaN $3.93\times 10^{-6}\%$
$y$ NaN $7.61\times 10^{-5}\%$ NaN $9.40\times 10^{-4}\%$
$z$ NaN $1.96\times 10^{-4}\%$ NaN $0.14\%$
Figure 1: Solutions to the Robertson ODEs computed using the numerical scheme selected by each LLM. For reference, the ODE system is also solved in scipy using the Radau method. For better visualization, the concentration for the $y$ species is scaled by $10^{4}$.
It is worth noting that both reasoning models, DeepSeek R1 and ChatGPT o3-mini-high, recognized the stiffness of the ODE system and chose an implicit method, while both non-reasoning models failed. During the reasoning, both DeepSeek R1 and ChatGPT o3-mini-high also gave explanations for their choice of implicit methods over explicit methods. DeepSeek R1 reported that
@@
Similarly, ChatGPT o3-mini-high reported that
@@
This experiment demonstrates the superior performance of the reasoning LLMs over non-reasoning ones, as reasoning models are able to analyze the problem first before choosing an appropriate method to implement.
@@ 2.1.2 Finite difference for Poisson equation
In this section, we test the ability of ChatGPT and DeepSeek models to implement finite difference methods for partial differential equations. In this experiment, we consider the following Poisson equation [44]
$\left\{\begin{array}[]{l}-\Delta u=f,x\in\Omega\\ u|_{\partial\Omega}=0,\end{array}\right.$
and set $\Omega=[-1,1]^{2}/[0,1]^{2}$ to be the L-shaped domain. The right hand side is set to be $f(x,y)=1$. The following question was asked to all four LLMs:
@@
To ensure fairness across different models, the question was asked in a new chat for each. Except for ChatGPT, all three other models provided error-free, executable codes. Notably, DeepSeek employed the most advanced iterative methods to solve the problem, whereas the others relied on conventional finite difference methods for discretization and solving the linear system. We plot the predictions of solution in Figure 2.
@@
We can clearly see that DeepSeek-R1 not only mistakes the sign of the solution but gives wrong scales, leading to worse performance. However, other models only mistake the sign of the solution, which means that the coding ability of these models surpasses the mathematical ability. Furthermore, for the comparison between the two reasoning models, ChatGPT o3-mini-high responds much faster than DeepSeek R1, with less thinking time.
Additional details are listed in Table 2. DeepSeek V3 takes much more time to give the results by using iteration methods. Overall, the performances of all models are comparable, but ChatGPT o3-mini-high can not only give accurate results but can save a lot of thinking time, leading to better performance.
Table 2: Summary of performance using different models.
DeepSeek V3 ChatGPT 4o DeepSeek R1 ChatGPT o3-mini-high
Method Iteration FDM FDM FDM
Grid $50\times 50$ $50\times 50$ $50\times 50$ $41\times 41$
CPU Time $9.5s$ $0.5s$ $0.6s$ $0.5s$
$L_{2}$ error 200% 200% 101% 200%
@@ 2.1.3 Finite element methods for Vibrating Beam equation
Here, $EI=1N/m^{2}$ is the constant flexural rigidity. For simplicity, we set $L=1$. The true solution is obtained by solving the ODE and given as
@@
Now, we ask all four LLMs the same question:
@@
Our results demonstrate that none of the four LLMs can correctly apply finite element methods to get the right solution. In detail, the predicted solutions obtained by these four models are demonstrated in Figure 5. Additional details can be found in Table 3.
Table 3: Summary of performance using different models.
DeepSeek V3 ChatGPT 4o DeepSeek R1 ChatGPT o3-mini-high
Basis Hermite Hermite Hermite Hermite
Grid 100 50 10 100
$L_{2}$ error 276300% 205% 153% 19.41%
We observe that all four LLMs use Hermite basis functions to approximate the solution. However, the non-reasoning models perform worse, showing significantly larger $L_{2}$ errors. In contrast, consistent with results from other experiments, the reasoning models solve the problem more effectively, with DeepSeek R1 achieving the lowest $L_{2}$ error.
@@ 2.1.4 Quadratures for integrals
The question posed to all four models is
@@
The approaches of different models vary significantly. DeepSeek-V3 directly uses the scipy.integrate module to compute the integral, while the other three first apply a transformation to remove the singularity. Specifically, by setting $x=t^{3}$, we can obtain
@@
After this transformation, the LLMs all chose Gaussian quadrature rules to approximate the integral. The additional details for the chosen methods and errors are summarized in Table 4.
Table 4: Summary of performance using different models.
DeepSeek V3 ChatGPT 4o DeepSeek R1 ChatGPT o3-mini-high
Method Scipy Gauss-Legendre Gauss-Legendre Gauss-Legendre
Number of nodes NAN 20 10 50
$L_{2}$ error NAN 5.15e-13 3.74e-12 5.21e-14
Overall, except DeepSeek-V3, all models achieve comparable performance. They are smart to make the transformation first and then apply quadrature rules. It can still be seen that the reasoning model ChatGPT o3-mini-high achieves the most accurate results using more nodes than the others.

### v2

# DeepSeek vs. ChatGPT vs. Claude: A Comparative Study for Scientific Computing and Scientific Machine Learning Tasks

@@ Abstract
Large Language Models (LLMs) have emerged as powerful tools for tackling a wide range of problems, including those in scientific computing, particularly in solving partial differential equations (PDEs). However, different models exhibit distinct strengths and preferences, resulting in varying levels of performance. In this paper, we compare the capabilities of the most advanced LLMs—DeepSeek, ChatGPT, and Claude—along with their reasoning-optimized versions in addressing computational challenges. Specifically, we evaluate their proficiency in solving traditional numerical problems in scientific computing as well as leveraging scientific machine learning techniques for PDE-based problems. We designed all our experiments so that a non-trivial decision is required, e.g. defining the proper space of input functions for neural operator learning. Our findings show that reasoning and hybrid-reasoning models consistently and significantly outperform non-reasoning ones in solving challenging problems, with ChatGPT o3-mini-high generally offering the fastest reasoning speed.
@@ 1 Introduction
Large language models (LLMs), such as ChatGPT [1, 2, 3], Claude [4], and, most recently, DeepSeek [5] are general-purpose artificial intelligence (AI) chatbots based on the transformer architecture [6], trained on large amounts of data from the Internet, and have learned to engage in conversations based on user requests. Among their many applications, there has been growing interest in leveraging LLMs for scientific research, particularly in coding, mathematics, and problem-solving. This interest has driven rapid recent advancements in LLM development. For example, on January 31, 2025, OpenAI released ChatGPT o3-mini [2], the latest reasoning model supposed to deliver particular strength in science, math, and coding. This came only a few days after DeepSeek unveiled its own reasoning model DeepSeek-R1 [7] that also specializes in math, coding, and reasoning tasks. Most recently, on February 24, 2025, Anthropic also announced Claude 3.7 Sonnet with extended thinking mode to improve its performance on math, physics, instruction-following, and coding [4]. These models build on prior advancements such as ChatGPT-4o [1], DeepSeek V3 [5], and Claude 3.5 Sonnet [8] marking a new phase of competition among LLM developers seeking to optimize performance in scientific fields.
@@
With the intensifying competition among LLM developers, this study conducts a comparative testing of their models in computational mathematics and scientific machine learning. Deliberately tricky problems are formulated and posed to both reasoning and non-reasoning models from the DeepSeek, ChatGPT, and Claude LLM families. We first test the LLMs’ capability in numerical methods, including numerical integration, finite difference methods (FDM), and finite element methods (FEM). Since machine learning methods have gained popularity as alternatives to traditional numerical approaches [38, 39], we also assess the LLMs’ performance in scientific machine learning tasks, including image recognition, physics-informed learning with physics-informed neural networks (PINNs) [40], and operator learning methods such as Deep Operator Network (DeepONet) [41]. By comparing the models’ performances and their reasoning decisions across a variety of benchmark problems, the work provides insight into how well these popular LLMs perform in computational mathematics and the risks and benefits of adopting them in our studies and research.
@@ 2 Experiments
In this section, we evaluate six different LLMs developed by DeepSeek, OpenAI, and Anthropic by testing them on a range of challenging benchmark problems in numerical algorithms and scientific machine learning. The models under consideration are:
@@
    ChatGPT 4o [1]: OpenAI’s current flagship model designed for general reasoning and multimodal capabilities, including text, code, and vision processing. It employs an optimized transformer architecture and focuses on improving efficiency and reducing latency while maintaining performance in coding, mathematics, and logical inference.
@@
    ChatGPT o3-mini-high [2]: ChatGPT o3-mini-high is a variant of OpenAI’s o3-mini model, optimized for tasks requiring intensive reasoning in coding, mathematics, and science. “High” stands for higher-intelligence. Despite its smaller size, o3-mini-high is supposed to deliver performance comparable to larger models by employing advanced reasoning techniques.
  5. 5.
    Claude 3.7 Sonnet [4]: Anthropic’s high-performance model designed for versatile text processing, code generation, and analytical tasks. It employs an optimized transformer architecture with enhanced contextual understanding capabilities and is supposed to deliver reliable performance across various domains including creative writing, programming, and knowledge-intensive applications. It claims to be the first “hybrid reasoning model” on the market.
  6. 6.
    Claude 3.7 Sonnet extended thinking [4]: An enhanced variant of Anthropic’s Sonnet model featuring an additional reasoning layer that enables deeper analytical thinking. This extended version incorporates a specialized processing mode that allows for more thorough evaluation of complex problems to improve performance on tasks requiring multi-step reasoning, mathematical problem-solving, and logical deduction while maintaining the base model’s versatility.
To ensure fairness, all chat memory and user-personalized settings are disabled. This prevents models from benefiting from prior context and ensures that each query is treated independently. Additionally, since LLM response times can be affected by server latency rather than actual evaluation speed, we do not compare or report response times for non-reasoning models (including the hybrid reasoning Claude 3.7 Sonnet). Instead, we only present in this paper the self-reported reasoning times for reasoning LLMs. The focus of our evaluation is on the quality of the generated solutions, and, for reasoning models, their reasoning process and decision-making. To differentiate the model performances, the test problems selected are tricky but also leave enough flexibility for the LLMs to make their own decisions in what method and parameters to use in their solutions. Most problems are advanced, typically at the PhD level, and demand a deep understanding of computational mathematics and familiarity with recent research advancements. The coding languages tested are chosen to be TensorFlow [42] and Jax [43] in Python, which are frameworks frequently used in the scientific machine learning community.
@@ 2.1 Traditional numerical methods
In this section, we present the six LLMs under consideration addressing a variety of problems in traditional numerical methods and solving differential equations. We compare the performance of their generated solutions and discuss the decisions the LLMs made.
@@ 2.1.1 Numerical solution to stiff ODEs
The Robertson problem Equation 1 is a well-known example of a stiff system of ODEs that describes the chemical reaction of H. H. Robertson [44]:
@@
The LLMs’ responses are summarized in Table 1, and the computed solutions are plotted in Figure 1. For reference, the ODE system is also solved in scipy using an implicit Runge-Kutta method of Radau IIA family of order 5 (Radau) [45].
@@
Model
Reasoning
time (s)
Method Step Size $L_{2}$ Error ($x$) $L_{2}$ Error ($y$) $L_{2}$ Error ($z$)
DeepSeek V3 N/A RK4 0.01 N/A N/A N/A
DeepSeek R1 249.0 Backward Euler 0.1 $7.86\times 10^{-6}\%$ $7.61\times 10^{-5}\%$ $1.96\times 10^{-4}\%$
ChatGPT 4o N/A RK4 0.1 N/A N/A N/A
ChatGPT 88.0 Backward Euler with Adaptive $3.93\times 10^{-6}\%$ $9.40\times 10^{-4}\%$ $0.14\%$
o3-mini-high adaptive time stepping ($10^{-6}$ initially)
Claude 3.7 Sonnet N/A RK4 with Adaptive $\mathbf{1.24\times 10^{-6}\%}$ $0.05\%$ $4.89\times 10^{-4}\%$
adaptive time stepping ($10^{-4}$ initially)
Claude 3.7 Sonnet
extended thinking
115.0 Backward Euler $0.1$ $7.74\times 10^{-6}\%$ $\mathbf{7.53\times 10^{-5}\%}$ $\mathbf{1.94\times 10^{-4}}\%$
Figure 1: Solutions to the Robertson ODEs computed using the numerical scheme selected by each LLM. For reference, the ODE system is also solved in scipy using the Radau method. For better visualization, the concentration for the $y$ species is scaled by $10^{4}$ in plots where the solution converges.
All non-reasoning models implemented a fourth-order Runge-Kutta (RK4) scheme. The versions by DeepSeek V3 and ChatGPT 4o chose relatively large step sizes, and they failed to solve the problem and yielded exponentially growing solutions. In contrast, the hybrid-reasoning Claude 3.7 Sonnet implemented an RK4 scheme with adaptive time stepping that compares one full step of size $\Delta t$ with two half‑steps (each of size $\Delta t/2$) to estimate the local truncation error. If the estimated error is below a prescribed tolerance, the step is accepted; otherwise, it is rejected and a new (smaller) time step is computed. Claude 3.7 Sonnet’s adaptive method yielded much better accuracy than the other two non-reasoning models, although errors on the $y$-species are still visible. On the other hand, all fully reasoning models implemented the backward Euler method. In particular, ChatGPT o3-mini-high used a similar adaptive time stepping method. Overall, Claude 3.7 Sonnet with extended thinking achieved the best results. ChatGPT o3-mini-high yielded satisfactory solutions with the shortest reasoning time.
We remark that all three reasoning models—DeepSeek R1, ChatGPT o3-mini-high, and Claude 3.7 Sonnet with extended thinking—recognized the stiffness of the ODE system and chose an implicit method, while all non-reasoning models failed to do so. During the reasoning process, DeepSeek R1 reported that
@@
ChatGPT o3-mini-high reported that
@@
Similarly, Claude 3.7 Sonnet with extended thinking reported that
"This is a stiff system of ODEs, characterized by the vastly different coefficients …I’ll
implement the Backward Differentiation Formula (BDF) method, which is specifically designed for
stiff problems."
This experiment demonstrates the superior performance of the reasoning LLMs over non-reasoning ones, as reasoning models are able to analyze the problem first before choosing an appropriate method to implement, resembling the approach of a human scientist.
@@ 2.1.2 Finite difference for Poisson equation
In this experiment, we test the capability of DeepSeek, ChatGPT, and Claude models to implement finite difference methods for partial differential equations. We consider the following Poisson equation [46]:
$\left\{\begin{array}[]{l}-\Delta u=f,\quad x\in\Omega\\ u|_{\partial\Omega}=0,\end{array}\right.$
and set $\Omega=[-1,1]^{2}/[0,1]^{2}$ to be an L-shaped domain. The right hand side function is set to be $f(x,y)=1$. We are specifically interested in testing if the LLMs can solve equations on a non-standard domain. The following question was asked to all of LLMs:
@@
Except for ChatGPT, all of other models provided error-free, executable codes. Notably, DeepSeek employed the most advanced iterative methods to solve the problem, whereas the others relied on conventional finite difference methods for discretization and solving the linear system. The predicted solutions are plotted in Figure 2.
[…]
@@
Results and methods used by each LLM are listed in Table 2. Among all six models, the only correct result is given by Claude 3.7 Sonnet. DeepSeek V3’s implementation takes a much longer CPU time to give the results due to its using iteration methods. DeepSeek-R1 not only mistakes the sign of the solution but gives wrong scales, leading to worse performance. All other models, except the correct Claude 3.7 Sonnet, only mistake the sign of the solution. Furthermore, among the three reasoning models, ChatGPT o3-mini-high responded much faster to requests than DeepSeek R1 and Claude 3.7 Sonnet with extended thinking.
Table 2: Summary of performance of different models in solving the Poisson equation.
Reasoning time(s) Method Grid CPU time (s) $L_{2}$ error
DeepSeek V3 N/A Iteration $50\times 50$ 9.5 201%
DeepSeek R1 476.0 FDM $50\times 50$ 0.6 203%
ChatGPT 4o N/A FDM $50\times 50$ 0.6 203%
ChatGPT o3-mini-high 32.0 FDM $41\times 41$ 0.5 207%
Claude 3.7 Sonnet N/A FDM $50\times 50$ 0.6 4.54%
Claude 3.7 Sonnet extended thinking 125.0 FDM $60\times 60$ 0.7 202%
@@ 2.1.3 Finite element methods for Vibrating Beam equation
Here, $EI=1N/m^{2}$ is the constant flexural rigidity. For simplicity, we set $L=1$. The true solution, plotted in Figure 3, is obtained by solving the ODE and given as
@@
All LLMs under consideration are asked with the following question:
@@
Results and details of the implementation can be found in Table 3. The solutions predicted by each LLM are plotted in Figure 6. Our results demonstrate that none of the LLMs can correctly apply finite element methods to get a very accurate solution.
Table 3: Summary of performance of different models in solving the beam equation.
Reasoning time (s) Basis Grid $L_{2}$ error
DeepSeek V3 N/A Hermite 100 276300%
DeepSeek R1 790.0 Hermite 10 15.3%
ChatGPT 4o N/A Hermite 50 205%
ChatGPT o3-mini-high 31.0 Hermite 100 19.4%
Claude 3.7 Sonnet N/A Hermite 100 957%
Claude 3.7 Sonnet extended thinking 9.0 Hermite 20 13.7%
We observe that all LLMs use Hermite basis functions to approximate the solution. Besides, they realize that the weak form should be derived and then the Galerkin methods can be applied. However, they all choose the wrong test space to derive the weak form, which leads to the final wrong results. Among these models, the non-reasoning models perform much worse than reasoning models, showing significantly larger $L_{2}$ errors. In contrast, consistent with results from other experiments, the reasoning models solve the problem more effectively, with Claude 3.7 Sonnet extended thinking achieving the smallest $L_{2}$ error. For response times among reasoning models, ChatGPT o3-mini-high and Claude 3.7 Sonnet with extended thinking responded much faster than DeepSeek R1.
Moreover, the implementations by Claude will further test convergence rates of the method using different numbers of basis functions. It turns out that Claude Sonnet models reason more technically like human scientists for mathematical problems and give comprehensive results to verify which method is better.
[…]
@@ 2.1.4 Quadratures for integrals
The question posed to all of the models is
@@
All LLMs provided error-free implementation, while the approaches of different models vary significantly. DeepSeek V3 directly used the scipy.integrate module to compute the integral, while the other models first applied a transformation to remove the singularity. Specifically, setting $x=t^{3}$ gives
@@
After applying this transformation, most LLMs employed Gaussian quadrature rules to approximate the integral. In particular, Claude models investigated a broader set of integration techniques to assess convergence rates, including the trapezoidal rule, Simpson’s rule, direct Gaussian quadrature, and the aforementioned variable transformation. As shown in Figure 5, the variable transform achieves the fastest convergence among these approaches, providing a thorough comparison of their performance.
Furthermore, in the case of Claude 3.7 Sonnet with extended reasoning, several quadrature methods are tested following the variable transformation—namely Gaussian quadrature, adaptive Simpson’s rule, and Romberg integration—offering additional insights into the efficacy of different numerical integration strategies. A detailed summary of the chosen methods and their associated errors can be found in Table 4.
Figure 5: Convergence rate of different methods tested for integration.
Table 4: Summary of performance of different models in integration.
Reasoning time (s) Method Number of Nodes $L_{2}$ error
DeepSeek V3 N/A Scipy N/A N/A
DeepSeek R1 584.0 Gauss-Legendre 10 3.74e-12
ChatGPT 4o N/A Gauss-Legendre 20 5.15e-13
ChatGPT o3-mini-high 74.0 Gauss-Legendre 50 5.21e-14
Claude 3.7 Sonnet N/A Gauss-Legendre 1000 Machine precision
Claude 3.7 Sonnet extended thinking 37.0 Gauss-Legendre 1000 Machine precision
Overall, except for DeepSeek V3, all models can achieve comparably high performance and correctly follow the user request. They are able to identify the singularity and correctly apply the transformation before using quadrature rules. Claude models provide more comprehensive comparisons of different methods and are more sophisticated in analyzing such scientific computing tasks.